OpenStudy (anonymous):

a survey was conducted among passengers who were travelling in bus from chennai to mumbai.it was found that 19 know tamil,17 know telugu,13 know hindi,11know two languages ane three languages. 5 know all the three languages.how many people are there in the bus?

OpenStudy (anonymous):

17-5=12 19-5=14 13-5=8 5+8+14+12=39 not saying its right but I ampretty sure of my answer

OpenStudy (anonymous):

thank you ..but this is not a correct ans.... the options are a) 29 b)43 c)31 d)30

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

11 knows two languages ane three languages..?